# Data-driven Error Estimation: Upper Bounding Multiple Errors without Class Complexity as Input

### v3

@@ Abstract
Constructing confidence intervals that are simultaneously valid across a class of estimates is central for tasks such as multiple mean estimation, bounding generalization error in machine learning, and adaptive experimental design. We frame this as an “error estimation problem,” where the goal is to determine a high-probability upper bound on the maximum error for a class of estimates. We propose an entirely data-driven approach that derives such bounds for both finite and infinite class settings, naturally adapting to a potentially unknown correlation structure of random errors. Notably, our method does not require class complexity as an input, overcoming a major limitation of existing approaches such as union bounding and bounds based on Talagrand’s inequality. In this paper, we present our simple yet general solution and demonstrate its flexibility through applications ranging from constructing multiple simultaneously valid confidence intervals to optimizing exploration in contextual bandit algorithms.
Machine Learning, ICML, Bandits, Generalization Bounds
@@ 1 Introduction
In statistical estimation, quantifying uncertainty involves bounding the error ($e$) between a target estimand ($\theta$) and its estimate ($\hat{\theta}$). While single errors are often bounded using concentration inequalities or the central limit theorem (CLT), many practical applications also require constructing simultaneous confidence intervals (CI) over multiple estimates. We frame this as the “error estimation problem”:
@@
For each $h$ in a class of estimation tasks $\mathcal{H}$, we define an estimand $\theta_{h}$, and its estimate $\hat{\theta}_{h}$. The error $e_{h}$, is a measure of distance between the estimand and the estimate. Our goal is to derive a high-probability upper bound on $\max_{h\in\mathcal{H}}e_{h}$.
@@
Consider for instance an educational assessment where $\mathcal{H}$ represents a (finite or infinite) set of various demographic or academic subgroups (such as ‘all students”,“male students”, “4th graders”, etc.). Setting $\theta_{h}$ to be the true mean test score for subgroup $h$, and $\hat{\theta}_{h}$ as its sample mean estimate, solving the error estimation problem would allow us to construct simultaneously valid CI for all subgroup mean test scores.
For finite estimation classes, a standard solution is the union bound, which adjusts individual error bounds based on the number of subgroups, $|\mathcal{H}|$. However, this method yields conservative bounds and doesn’t extend to cases with infinitely many subgroups. Another example of an infinite class arises when bounding the excess risk of a model – the gap between a learned model’s error and the error of the best model in the model class – can be framed as an error estimation problem over the model class. Existing methods for infinite classes rely on the need to characterize class complexity (e.g.VC dimension or Rademacher complexity). These measures can be hard to determine precisely, leading to the use of upper bounds which result in overly conservative CI in practice.
Motivated by these limitations, in this paper we present a new data-driven method for constructing valid confidence intervals. Our approach: (a) circumvents the need for class complexity characterization - a major limitation that, to the best of our knowledge, has not been overcome before, and (b) naturally adapts to the true correlation structure of errors, yielding much tighter bounds in cases where traditional methods are overly conservative. We achieve this goal by leveraging single error bounds, which can be either finite-sample or asymptotic, to create corresponding uniform error bounds through a surprisingly simple procedure (Theorem 2.2).
To provide intuition for our approach, we draw an analogy between our solution to bounding maximum error over a class and the well-known overfitting phenomenon that often occurs in machine learning. The goal in many learning problems is to minimize true risk (expected loss over a data distribution) over a class of predictors $\mathcal{F}$. Empirical risk minimization (ERM) typically approximates this by selecting $f\in\mathcal{F}$ to minimize empirical risk (average loss) on some dataset $S$. This inherently “overfits” to $S$: with high probability, the minimum empirical risk is less than the minimum true risk (up to a small correction). [^1] This is closely related to the fact that training error is nearly always less than test error. Formally, if we let $\mathcal{R}(f)$ be the true risk and ^$\hat{\mathcal{R}}(f)$ be the empirical risk, we see with high probability
The rest of the paper is organized as follows: Section 1.1 provides a comprehensive review of related literature. Section 2 presents the core theoretical framework for error estimation, structured into three key subsections.
Section 2.1 introduces the fundamental setting and notation. Section 2.2 presents the main idea in our approach. Building upon this foundation, Section 2.3 proposes an advanced localization procedure to refine the bounds derived in Section 2.2.
Section 3 then demonstrates how to apply the error estimation method in practice. We first outline a general recipe in Section 3.1, followed by two examples. In Section 3.2, we apply the error estimation theory derived in Section 2 to bound the maximum error in the multiple mean setting. Section 3.3 illustrates how to extend the theory to bound excess risk in supervised learning settings. We conclude by discussing the broader implications of our work.
@@
Additionally, in Appendix C we provide an extension of the error estimation model to multiple hypothesis testing. Furthermore, Appendix D presents a key application of our work to contextual bandits, a challenging problem in adaptive experimentation. Contextual bandit algorithms require a measure of estimation error/uncertainty to guide exploration. While existing algorithms (with tight theoretical guarantees) rely on mathematically derived class complexity bounds to quantify such errors (Foster & Rakhlin, 2023), our error estimation approach can achieve tight error/uncertainty bounds without needing such a priori mathematical analysis. [^2] For many classes of interest, such as neural networks, we don’t have a precise mathematical characterization of class complexity and this is still an active area of research. Building on insights from (Foster et al., 2020; Krishnamurthy et al., 2023, 2024), we show how error estimation enables the development of theoretically sound contextual bandit algorithms that accommodate a wide range of policy classes.
@@ 1.1 Related Work
Error estimation is closely tied to generalization bounds in statistical learning theory. These bounds quantify the gap between a model’s empirical performance on training data and its expected performance on unseen data, uniformly across all models in a specified class. Such bounds typically rely on measures or bounds on the complexity of the model class (Shalev-Shwartz & Ben-David, 2014; Koltchinskii, 2011).
@@
Bootstrapping is a popular empirical alternative for constructing confidence intervals, particularly effective in the case of large but finite classes, such as when estimating the mean of a high-dimensional random variable. When statistical complexity is bounded by a polynomial function of sample size, bootstrap-based methods have been applied to construct asymptotically valid confidence intervals around the maximum ($L^{\infty}$) error (see Giné & Zinn, 1990; Chernozhukov et al., 2013, 2023). However, bootstrapping faces several limitations: (1) it is effective in high but not infinite dimensions, limiting its applicability in continuous class settings, (2) it only provides asymptotic guarantees, and (3) the number of bootstrap samples, $B$ must be chosen heuristically and scales poorly with model complexity. This becomes computationally expensive as bootstrapping requires solving $B$ optimization problems. Our approach, on the other hand, simplifies to solving a single optimization problem.
The literature on PAC-Bayes bounds (see Alquier, 2021, for a user friendly tutorial) is also motivationally related to error estimation. PAC-Bayes seeks to improve upon class complexity-based bounds in traditional statistical learning theory by leveraging a prior over the given hypothesis class. With an informed prior that has a high-probability of sampling hypotheses with low average loss, the resulting generalization bounds can be much better than traditional class complexity-based bounds. However, when the prior is uninformative (e.g. uniform-like distribution over hypotheses), PAC-Bayes bounds have no advantage over traditional class complexity-based bounds. Thus, two key disadvantages with PAC-Bayes bounds are (1) the need to work with a distribution over hypotheses (ERM only outputs a single hypothesis and does not fall into this framework) and (2) a need to specify good priors. Our solution to error estimation simultaneously bounds every error of interest, and doesn’t require an informative prior to achieve an improvement over traditional methods.
The statistical learning theory bounds described above are critical for theoretically sound algorithms for active learning, contextual bandits, and reinforcement learning (RL) (Agarwal et al., 2019; Foster & Rakhlin, 2023). For contextual bandits, we replace the need to know class complexity-based bounds with a data-driven estimation of the error. One can imagine that error estimation can similarly improve on the statistical learning theory bounds in active learning and RL. Other potential areas of applications include: detecting model degradation from tracking a class of evaluation tasks, multiple hypothesis testing in genetic analysis, etc.
@@
It is helpful to note that our work isn’t closely related to literature on constructing anytime-valid confidence intervals (CIs) (e.g., Ramdas et al., 2023). While both error estimation and the literature on anytime-valid CIs involve constructing simultaneously valid CIs, we do not explicitly model or consider time dependent structure in our data.
@@ 2 Error Estimation Theory
2.1 Preliminaries for Error Estimation
We begin by formally defining the error estimation problem. Consider an initial random variable $S_{\text{def}}$ representing a dataset. This will be the “defining” dataset, used to construct estimates and define errors. Now we introduce a set or class of indices $\mathcal{H}$ (the set of subgroups in the educational assessment example), where each element $h\in\mathcal{H}$ is an index to an estimation task with three associated quantities: an unknown target estimand $\theta_{h}$ (the true mean test scores for subgroup $h$), a given estimate $\hat{\theta}_{h}$ estimated from $S_{\text{def}}$ (the sample mean test scores), and an unknown real-valued error $e_{h}$. Here the error ($e_{h}$) is an unknown real-valued measure of “proximity” (often a normalized distance) between the given estimate ($\hat{\theta}_{h}$) and the unknown estimand ($\theta_{h}$). Let $\mathcal{E}$ be a function that given a realization $s$ of $S_{\text{def}}$, produces the class of estimand-estimate-error tuples associated with each estimation task/index. $\mathcal{E}(s):=\{(h,\hat{\theta}_{h}(s),\theta_{h},e_{h}(s))|h\in\mathcal{H}\}$, where $\hat{\theta}_{h}(s)$ is the estimate for task $h$ computed from dataset realization $s$, and $e_{h}(s)$ is the corresponding error. [^3] In fact, all quantities in $\mathcal{E}$ can depend on s, see Section 3.3, Appendix C, and Section D.2 for examples where $\mathcal{H}$ and/or $\theta_{h}$ depend on $s$ as well. To lighten the notation, we refer to $e_{h}(s)$ and $\hat{\theta}_{h}(s)$ as simply $e_{h}$ and $\hat{\theta}_{h}$ where the dependence on $s$ is clear from context.
@@ 2.1 Preliminaries for Error Estimation
Now consider another random variable $S_{\text{err}}$, representing the “error estimation dataset”. This dataset is drawn from some distribution $\mathcal{D}_{s}$ and will be used to construct bounds on the maximum error $\max_{h\in\mathcal{H}}e_{h}(s)$. Here, $\mathcal{D}_{s}$ may be parameterized by quantities depending on the realization $s$ of $S_{\text{def}}$. [^4] For example, in reinforcement learning, future sampling distributions may depend on earlier estimates. However, in other estimation settings, estimating some parameters of a given distribution would not affect the distribution of future/holdout samples. Once $\mathcal{D}_{s}$ is defined, $S_{\text{err}}$ and its realizations do not further depend on $s$. We often enforce such conditions by ensuring that the random dataset $S_{\text{err}}$ is either a newly sampled dataset from $\mathcal{D}_{s}$ or a holdout dataset that wasn’t used to construct any of the estimates $\hat{\theta}_{h}(s)$ in $\mathcal{E}(s)$.
For any given confidence parameter $\delta\in(0,1)$, the goal of error estimation is thus to use the random dataset $S_{\text{err}}$ to estimate a high-probability (holds with probability at least $1-\delta$) upper bound on the maximum error ($\max_{h\in\mathcal{H}}e_{h}$).
Like the union bound, our solution to the error estimation problem assumes access to an estimator ($\hat{u}:(S_{\text{err}},h,\lambda)\to\mathbb{R}$) that outputs, for any single $h\in\mathcal{H}$, an upper bound on the error $e_{h}$ with probability at least $1-\lambda$ (under the distribution of $S_{\text{err}}$) (see 2.1). Our main result (Theorem 2.2) then leverages $\hat{u}$ in order to construct a high-probability upper bound on the maximum error (that is, $\max_{h\in\mathcal{H}}e_{h}$) – without needing bounds on the complexity or size of the class $\mathcal{H}$ as an input.
@@
We have access to error upper bound estimators $\hat{u}:(S_{\text{err}}\times\mathcal{H}\times(0,1))\to\mathbb{R}$ that satisfy the following condition:
$\Pr_{S_{\text{err}}\sim\mathcal{D}}(\hat{u}(S_{\text{err}},h,\lambda)\geq e_{h}|\mathcal{E}(s))\geq 1-\lambda$
for all $h\in\mathcal{H}$ and $\lambda\in(0,1)$.
2.2 Main Result: Data-Driven Error Estimation
We now describe our surprisingly simple and hence very general solution (Theorem 2.2) to the error estimation problem, which notably doesn’t require any explicit information on the size/complexity of the class $\mathcal{H}$ [^5] The validity of our solution doesn’t require any condition on the class $\mathcal{H}$ to hold. For instance, it accommodates any measure of complexity of $\mathcal{H}$ to be infinitely large.
@@ 2.2 Main Result: Data-Driven Error Estimation
Suppose 2.1 holds. Let $\hat{\xi}$ denote our estimated upper bound on the maximum error, defined in (2). Then for any $\delta\in(0,1)$, we have that (3) and (4) hold. [^6] Neither (2) (definition of $\hat{\xi}$) or (3) (main guarantee) explicitly depend on the complexity size of class $\mathcal{H}$.
$\hat{\xi}(S_{\text{err}},\delta):=\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ (2)
@@
Consider an index achieving maximum error, $\tilde{h}\in\arg\max_{h\in\mathcal{H}}e_{h}$ and event $\mathcal{W}(S):=\{\hat{u}(S,\tilde{h},\delta)\geq e_{\tilde{h}}\}$. We first show that $\mathcal{W}(S_{\text{err}})$ implies that $\hat{\xi}(S_{\text{err}},\delta)$ upper bounds the maximum error.
$\mathcal{W}(S_{\text{err}})$ $\implies e_{\tilde{h}}\stackrel{{\scriptstyle(i)}}{{\leq}}\hat{u}(S_{\text{err}},\tilde{h},\delta)\stackrel{{\scriptstyle(ii)}}{{\leq}}\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ (5)
$\stackrel{{\scriptstyle(iii)}}{{\implies}}\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}.$
Here (i) follows from $\mathcal{W}(S_{\text{err}})$, (ii) follows from $\tilde{h}\in\mathcal{H}$, and (iii) follows from (2) (definition of $\hat{\xi}$) and definition of $\tilde{h}$. We now utilize (5) and 2.1 to show that (3) holds.
$\Pr_{S{\text{err}}\sim\mathcal{D}}\Big{(}\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\Big{|}\mathcal{E}(s)\Big{)}\stackrel{{\scriptstyle(i)}}{{\geq}}\Pr_{S{\text{err}}\sim\mathcal{D}}(\mathcal{W}(S_{\text{err}})|\mathcal{E}(s))$ (6)
=$\stackrel{{\scriptstyle(ii)}}{{=}}\Pr_{S_{\text{err}}\sim\mathcal{D}}\Big{(}\hat{u}(S_{\text{err}},\tilde{h},\delta)\geq e_{\tilde{h}}\Big{|}\mathcal{E}(s)\Big{)}\stackrel{{\scriptstyle(iii)}}{{\geq}}1-\delta$
Here (i) follows from (5), (ii) follows from the definition of $\mathcal{W}$, and (iii) follows from 2.1 – showing that (3) holds. Now by the law of total expectation [^7] If we let $I=\mathbf{1}_{\{\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\}}$, the law of total expectation (or tower rule) states that $\mathbb{E}_{S_{\text{def}}}[\mathbb{E}_{S_{\text{err}}}[I|\mathcal{E}(s)]]=\mathbb{E}_{S_{\text{def}},S_{\text{err}}}[I]$ and (3), we also have that (4) holds. This completes the proof of Theorem 2.2.
Theorem 2.2 shows that, with probability at least $1-\delta$, $\hat{\xi}(S_{\text{err}},\delta)$ is an upper bound on max error ($\max_{h\in\mathcal{H}}e_{h}$). In contrast to a union bound approach, the definition of $\hat{\xi}(S_{\text{err}},\delta)$ does not explicitly depend on the complexity or size of class $\mathcal{H}$. This allows our method to naturally adapt to the correlation structure of the errors. As a consequence, our bound matches the union bound (up to constant factors) when it is tight, while providing much tighter bounds when it is loose (e.g., for highly correlated errors, see Remark 3.4 for a concrete example). Note that the dependence on $\mathcal{H}$ occurs in the maximization over the class $\mathcal{H}$. For finite classes, this maximization is straightforward, while for infinite classes it can be accomplished using standard optimization procedures such as gradient descent.

### v4

# Data-driven Error Estimation: Excess Risk Bounds without Class Complexity as Input

@@ Abstract
Constructing confidence intervals that are simultaneously valid across a class of estimates is central to tasks such as multiple mean estimation, generalization guarantees, and adaptive experimental design. We frame this as an “error estimation problem,” where the goal is to determine a high-probability upper bound on the maximum error for a class of estimates. We propose an entirely data-driven approach that derives such bounds for both finite and infinite class settings, naturally adapting to a potentially unknown correlation structure of random errors. Notably, our method does not require class complexity as an input, overcoming a major limitation of existing approaches. We present our simple yet general solution and demonstrate applications to simultaneous confidence intervals, excess-risk control and optimizing exploration in contextual bandit algorithms.
Machine Learning, ICML
@@ 1 Introduction
We introduce the following “error estimation problem”: For each $h$ in a class of estimation tasks $\mathcal{H}$, we define a true estimand $\theta_{h}$ and its estimate $\hat{\theta}_{h}$. The error $e_{h}$, is a measure of distance between the estimand and the estimate. Our goal is to derive a high-probability upper bound on $\max_{h\in\mathcal{H}}e_{h}$.
Example 1: Subgroup Means
@@
Consider an educational assessment where $\mathcal{H}$ represents a (finite or infinite) set of various demographic or academic subgroups (such as ‘all students”, “male 4th graders”, etc.). Setting $\theta_{h}$ to be the true mean test score for subgroup $h$, and $\hat{\theta}_{h}$ as its sample mean estimate, we may wish to construct simultaneously valid confidence intervals (CI) for all subgroup mean test scores. By providing an upper bound for $\max_{h\in\mathcal{H}}e_{h}$, a solution to the error estimation problem directly yields such intervals.
Example 2: Generalization and Excess Risk
A second example, with an infinite class of tasks, arises in supervised learning. The excess risk (Shalev-Shwartz and Ben-David, 2014) – the gap between a learned model’s error and the error of the best model in the model class – is a critical input into sequential decision-making tasks, such as contextual bandits, where the bounds are needed to inform the tradeoff between exploration (gathering more data in order to identify better policies) and exploitation (taking advantage of a policy that is near-optimal) (Agarwal et al., 2019). We approach this by setting up an error estimation problem where $\mathcal{H}$ represents a model class, $\theta_{h}$ is the true risk of a model $h\in\mathcal{H}$, and $\hat{\theta}_{h}$ is the empirical risk over a dataset. Bounding $\max_{h\in\mathcal{H}}e_{h}$ then gives us a generalization bound across $\mathcal{H}$, which can be combined with localization procedures to control excess risk.
The literature has proposed several approaches to the error estimation problem. While single errors are often bounded using concentration inequalities or the central limit theorem (CLT), the problem of constructing simultaneous guarantees over an entire family of estimates is more challenging. For finite estimation classes (e.g. a finite number of subgroups, as in Example 1), a standard (but conservative) solution is the union bound, which adjusts individual error bounds based on the number of subgroups, $|\mathcal{H}|$; practical empirical alternatives such as bootstrapping are widely used. Existing methods for infinite classes express uncertainty as a function of class complexity (e.g. VC dimension or Rademacher complexity), which is often difficult to characterize, leading to the use of conservative estimates of complexity and thus overly conservative upper bounds and resulting CIs. Nonetheless, no widely adopted empirical alternatives exist.
Contributions.
(i) A data-splitting wrapper for uniform guarantees (finite-sample under finite-sample pointwise bounds): given any pointwise high-probability error bound that holds for each fixed $h$, we obtain a uniform bound on $\sup_{h\in\mathcal{H}}e_{h}$. This applies to both finite and infinite classes and automatically adapts to correlations across tasks  [^1] Although these are uniform bounds, in our setup $e_{h}$ can be normalized for each $h$, allowing us to get simultaneously valid variance dependent bounds – see Section 3.2. (ii) A computable, data-dependent complexity proxy: in supervised learning, the maximization objective induced by our framework coincides with a standard symmetrization/discrepancy quantity and yields a practical analogue of Rademacher-style complexity (formal connections and tightness in Appendix A). (iii) Data-driven localization: when uniform error bounds are unnecessary, we use an iterative localization procedure that yields tighter bounds by progressively excluding models less relevant to the objective—for example, excluding high-risk models when bounding excess risk. (iv) Contextual bandits beyond realizability: we show how error estimation yields exploration sets and bonuses that can be plugged into modular contextual bandit pipelines while retaining theoretical guarantees.
@@
Contextual bandit algorithms require a measure of estimation error/uncertainty to guide exploration. While existing algorithms (with tight theoretical guarantees) rely on mathematically derived class complexity bounds to quantify such errors (Foster and Rakhlin, 2023), our error estimation approach can achieve tight error/uncertainty bounds without needing such a priori mathematical analysis. Building on insights from (Foster et al., 2020; Krishnamurthy et al., 2023, 2024), we show how error estimation enables the development of theoretically sound contextual bandit algorithms that accommodate a wide range of policy classes.
@@ 1.1 Related Work
Our approach to the error estimation problem can be contrasted with several alternatives. For the finite case, beyond the classical literature on multiple hypothesis testing (Efron, 2010), bootstrapping is particularly effective in the case of large but finite classes with potentially correlated errors (see e.g. Romano and Wolf, 2005). When statistical complexity is bounded by a polynomial function of sample size, bootstrap-based methods have been applied to construct asymptotically valid confidence intervals for the maximum ($L^{\infty}$) error (Giné and Zinn, 1990; Chernozhukov et al., 2013, 2023). However, (1) its guarantees typically require finite-dimensional or high-dimensional regimes with additional regularity/complexity control, (2) it only provides asymptotic guarantees, and (3) the number of bootstrap samples, $B$ must be chosen heuristically and scales poorly with model complexity.
In the infinite case, simultaneous error bounds are typically derived by analyzing Rademacher complexity or VC dimension (Shalev-Shwartz and Ben-David, 2014; Koltchinskii, 2011). These bounds are difficult to analyze and often fail to explain generalization behavior in modern overparameterized settings (Zhang et al., 2017). When our framework is applied to the generalization setting from Example 2, the resulting objective recovers an empirical analogue of Rademacher complexity, with data splitting playing the role of a symmetrization trick in Rademacher-based proofs. This coincides with the discrepancy function introduced in (Bartlett et al., 2002). However, our broader framework (i) holds under weaker distribution assumptions (allowing for applications to the contextual bandit setting), and (ii) is empirically validated in representative settings. We also show that the objective is upper-bounded by the true Rademacher complexity (see Appendix A), ensuring both practical feasibility and theoretical guarantees. Section 2.3 introduces a data-driven localization procedure, analogous to theoretical localized complexities (Koltchinskii, 2011), allowing us to tighten excess risk bounds with no additional mathematical analysis.
PAC-Bayes bounds (Alquier, 2021) offer an alternative to classical complexity-based generalization bounds by incorporating a prior over the hypothesis class (Dziugaite and Roy, 2017a). While good priors, concentrated around effective models, can provide tighter generalization bounds, uninformative priors result in no advantage over traditional class complexity-based bounds. Specifying good priors is hard in practice (Dziugaite and Roy, 2017b). Our solution to error estimation simultaneously bounds every error of interest, and doesn’t require an informative prior to achieve an improvement over traditional methods.
The statistical learning theory bounds described above are critical for theoretically sound algorithms for active learning, contextual bandits, and reinforcement learning (RL). For instance, exploration strategies typically rely on the complexity of the policy class as an explicit input to estimate uncertainty (Agarwal et al., 2014; Foster et al., 2020). Our framework replaces these bounds with empirically estimated quantities, enabling a data-driven estimation of the error. We anticipate that error estimation can similarly improve on the statistical learning theory bounds in active learning and RL.
Our work is also related in spirit to conformal prediction, which uses data splitting to construct marginal prediction intervals with coverage around individual test points (Vovk et al., 2005). While both approaches share the underlying idea of reserving a portion of the data to assess uncertainty, the goals and guarantees differ. Conformal methods provide marginal coverage for predictions on new, unseen instances, while our framework provides uniform error bounds over a class of estimated parameters.
@@
It is helpful to note that our work is not closely related to constructing anytime-valid CIs (e.g., Ramdas et al., 2023), as we do not explicitly model or consider time dependent structure in our data.
@@ 2 Error Estimation Theory
In this section, we introduce fundamental notation for error estimation (Section 2.1), present the main idea (Section 2.2) and propose a localization procedure for refined bounds (Section 2.3).
2.1 Problem Setup
We begin by formalizing the error estimation problem. We fix an index set $\mathcal{H}$ that labels a (finite or infinite) family of estimation tasks (the set of subgroups in Example 1), and a random defining dataset $S_{\text{def}}$. For each $h\in\mathcal{H}$, we associate three quantities.
  • $\theta_{h}$: an unknown target estimand (the true mean test score for subgroup $h$),
  • $\hat{\theta}_{h}$: an estimate of $\theta_{h}$ estimated from $S_{\text{def}}$ (sample mean test score for subgroup $h$),
  • $e_{h}$: an unknown real-valued error measuring “proximity” between $\hat{\theta}_{h}$ and $\theta_{h}$.
Let $\mathcal{E}$ be a function that maps a realization $s$ of $S_{\text{def}}$, to the class of estimand-estimate-error tuples associated with each estimation task/index. $\mathcal{E}(s):=\{(h,\hat{\theta}_{h}(s),\theta_{h},e_{h}(s))|h\in\mathcal{H}\}$. [^2] In some applications $\mathcal{H}$ or $\theta_{h}$ may themselves depend on $S_{\text{def}}$; see Sections 3.3, F, and H.2. To lighten the notation, we refer to $e_{h}(s)$ and $\hat{\theta}_{h}(s)$ as simply $e_{h}$ and $\hat{\theta}_{h}$ where the dependence on $s$ is clear from context. For a given confidence parameter $\delta\in(0,1)$, the goal of error estimation is to estimate a high-probability (holds with probability at least $1-\delta$) upper bound on the maximum error ($\max_{h\in\mathcal{H}}e_{h}$).
@@ 2.1 Problem Setup
To do so, we introduce $S_{\text{err}}$, an independent error estimation dataset, drawn from some distribution $\mathcal{D}$, that may depend on the realized $S_{\text{def}}=s$. [^3] For example, in reinforcement learning, future sampling distributions may depend on earlier estimates. However, in other estimation settings, estimating some parameters of a given distribution would not affect the distribution of future/holdout samples. Once $\mathcal{D}$ is defined, $S_{\text{err}}$ and its realizations do not further depend on $s$. We often enforce such conditions by ensuring that the random dataset $S_{\text{err}}$ is either a newly sampled dataset from $\mathcal{D}$ or a holdout dataset that wasn’t used to construct any of the estimates $\hat{\theta}_{h}(s)$ in $\mathcal{E}(s)$.
Bounding a single error $e_{h}$ is often straightforward using classical concentration inequalities or asymptotic results. Bounding the maximum, $\max_{h\in\mathcal{H}}e_{h}$, on the other hand, is more statistically challenging. A common approach is to build on single error bounds to construct a uniform guarantee. For instance, the union bound applies a correction to single error confidence levels to ensure simultaneous validity. Similarly, our solution to the error estimation problem assumes access to a single high-probability estimator (see 2.1). Our main insight is that this problem can be reduced to Theorem 2.2, which constructs a high-probability upper bound on the maximum error (that is, $\max_{h\in\mathcal{H}}e_{h}$) – without needing bounds on the complexity or size of the class $\mathcal{H}$ as an input.
@@
We have access to error upper bound estimators $\hat{u}:(S_{\text{err}}\times\mathcal{H}\times(0,1))\to\mathbb{R}$ that satisfy the following condition: for all $h\in\mathcal{H}$ and $\lambda\in(0,1)$,
$\Pr_{S_{\text{err}}\sim\mathcal{D}}(\hat{u}(S_{\text{err}},h,\lambda)\geq e_{h}|\mathcal{E}(s))\geq 1-\lambda.$
2.2 Main Insight: Data-Driven Error Estimation
We now formalize the key technical insight: Uniform control over $\max_{h\in\mathcal{H}}e_{h}$ reduces to constructing pointwise $(1-\delta)$ bound that holds for every fixed $h$. Since the maximizing index $\tilde{h}\in\arg\max_{h\in\mathcal{H}}e_{h}$ is measurable with respect to $S_{\text{def}}$, conditioning makes it fixed with respect to the randomness in $S_{\text{err}}$, yielding a uniform guarantee without an explicit complexity term.
@@ 2.2 Main Insight: Data-Driven Error Estimation
Suppose 2.1 holds. Let $\hat{\xi}$ denote our estimated upper bound on the maximum error, $\hat{\xi}(S_{\text{err}},\delta):=\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$. Then for any $\delta\in(0,1)$, we have that (1) and (2) hold. [^4] Neither the definition of $\hat{\xi}$ or (1) (main guarantee) explicitly depend on the complexity size of class $\mathcal{H}$. [^5] When $\mathcal{H}$ is an infinite class, replace $\max_{h\in\mathcal{H}}$ by $\sup_{h\in\mathcal{H}}$. See Appendix D for details
@@
Let the index $\tilde{h}\in\arg\max_{h\in\mathcal{H}}e_{h}$ and the event $\mathcal{W}(S):=\{\hat{u}(S,\tilde{h},\delta)\geq e_{\tilde{h}}\}$. We first show that $\mathcal{W}(S_{\text{err}})$ implies that $\hat{\xi}(S_{\text{err}},\delta)$ upper bounds the maximum error. Under $\mathcal{W}(S_{\text{err}})$, we have $e_{\tilde{h}}=\max_{h\in\mathcal{H}}e_{h}$ and
$e_{\tilde{h}}\stackrel{{\scriptstyle(i)}}{{\leq}}\hat{u}(S_{\text{err}},\tilde{h},\delta)\stackrel{{\scriptstyle(ii)}}{{\leq}}\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ $\stackrel{{\scriptstyle(iii)}}{{=:}}\hat{\xi}(S_{\text{err}},\delta).$ (3)
Here (i) follows from $\mathcal{W}(S_{\text{err}})$, (ii) follows from $\tilde{h}\in\mathcal{H}$, and (iii) follows from the definition of $\hat{\xi}$. We now utilize (3) and 2.1 to show that (1) holds.
$\Pr_{S{\text{err}}\sim\mathcal{D}}$ $\Big(\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\Big|\mathcal{E}(s)\Big)$ (4)
$\stackrel{{\scriptstyle(i)}}{{\geq}}\Pr_{S_{\text{err}}\sim\mathcal{D}}\Big(\hat{u}(S_{\text{err}},\tilde{h},\delta)\geq e_{\tilde{h}}\Big|\mathcal{E}(s)\Big)$ $\stackrel{{\scriptstyle(ii)}}{{\geq}}1-\delta$
Here (i) follows from (3) and the definition of $\mathcal{W}$, and (ii) follows from 2.1 – showing that (1) holds. Now by the law of total expectation [^6] If $I=\mathbf{1}_{\{\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\}}$, the law of total expectation states that $\mathbb{E}_{S_{\text{def}}}[\mathbb{E}_{S_{\text{err}}}[I|\mathcal{E}(s)]]=\mathbb{E}_{S_{\text{def}},S_{\text{err}}}[I]$ and (1), we also have that (2) holds. ∎
Theorem 2.2 is a general reduction: pointwise finite-sample error bounds on $S_{\text{err}}$ immediately yield uniform guarantees over $S_{\text{def}}$-dependent classes. In contrast to a union bound approach, the definition of $\hat{\xi}(S_{\text{err}},\delta)$ does not explicitly depend on the complexity or size of class $\mathcal{H}$. Note that the dependence on $\mathcal{H}$ occurs implicitly in the maximization over the class $\mathcal{H}$. For finite classes, this maximization is straightforward. For infinite classes, our guarantee is stated for $\sup_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$; in practice we approximate this supremum with optimization heuristics (e.g., SGD with multiple restarts).